OpenStudy (anonymous):

similar to FOIL: (sqrt3+sqrt5)^2

OpenStudy (johnweldon1993):

Right so \[\large (\sqrt{3} + \sqrt{5})^2 = (\sqrt{3} + \sqrt{5})(\sqrt{3} + \sqrt{5})\] And foil that out Hint \(\large \sqrt{a} \times \sqrt{b} = \sqrt{a \times b}\)

OpenStudy (anonymous):

its just so complicated! :( @johnweldon1993

OpenStudy (johnweldon1993):

No problem, alright, so you know how foil works right?

OpenStudy (anonymous):

its just distributing?

OpenStudy (johnweldon1993):

Basically yes, You just take the first term in the first parenthesis, and distribute it to the second parenthesis, then do the same with the second term.

OpenStudy (johnweldon1993):

So here, our first term in the first parenthesis is √3 right? Distribute that into the second parenthesis

OpenStudy (anonymous):

so sqrt3 into the other sqrt3?

OpenStudy (johnweldon1993):

Correct...so what is √3 times √3 ? Remember that hint i gave you above √a * √b = √(a * b)

OpenStudy (anonymous):

9

OpenStudy (johnweldon1993):

√9... \[\large \sqrt{3} \times \sqrt{3} = \sqrt{3 \times 3} = \sqrt{9}\] right? which we know to be \[\large \sqrt{9} = 3\]

OpenStudy (anonymous):

okay then what?

OpenStudy (johnweldon1993):

Well that was only distributing it to 1 thing in the parenthesis, we still need to distribute it to the √5

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

@johnweldon1993